# A Novel Approach to Industrial Defect Generation through Blended Latent Diffusion Model with Online Adaptation

### v1

@@ III-B3 Defect trimaps
As shown in Fig. 2, in the proposed AdaBLDM algorithm, the controlling trimap $\Gamma$ is firstly embedded by a convolutional block as $\zeta(\Gamma)\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$. The embedding feature is then fed into an encoder network $\hat{\epsilon}_{tri}(\cdot)$ that shares the same structure and initial parameters with the denoising encoder of $\epsilon_{\theta}(\cdot)$, except that the original “spatial attention” module is replaced with a self-attention process over the embedded trimap features. The deep features extracted from different layers of the denoising encoder are merged with the intermediate features of the denoising decoder of LDM via element-wise addition (see Fig. 2).
III-C Training Scheme
Formally, our cross-modal controlling features is defined as $\mathcal{C}=\{\tau(y),\Gamma\}$ and the learning objective of LDM is defined as:
$L_{LDM}\coloneqq\mathbb{E}_{\mathbf{z}_{t},\mathcal{C},\epsilon\thicksim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\mathcal{C})\|^{2}_{2}\right].$ (8)
In this work, the denoising model ${\epsilon}_{\theta}(\cdot)$, the embedding block $\zeta(\cdot)$ of trimap prompts and the encoder $\hat{\epsilon}_{tri}$ for the latent trimap feature are all learned via optimizing the loss function defined in Eq. 8. However, different training schemes are applied for different modules. In specific, ${\epsilon}_{\theta}(\cdot)$ is firstly pre-trained by using the domain dataset (e.g. all the training images of all the subcategories of MVTec AD [19] for the experiments on this dataset) and then frozen during model learning for each specific subcategory; the latent feature encoder $\hat{\epsilon}_{tri}(\cdot)$ is initialized by the corresponding parameters of the original LDM model [16] and then fine-tuned for each subcategory based on the real defective samples; the convolutional block $\zeta(\cdot)$ is only learned based on the genuine anomaly samples with a random initialization.
Apart from the reverse diffusion model, the text prompts $y$ are obtained by using the CLIP [48] and BLIP [49] algorithm for sufficiently various descriptions. The text encoder is copied from CLIP and frozen during training. As to the image encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$, we follow [16] to inherit the deep models of the VQ-VAE algorithm [50] and fix them during the whole training stage.
@@ III-D Multi-Stage Denoising with Content Editing
Inspired the seminal work, namely Blended Latent Diffusion (BLD) [17], we propose to perform feature editing in the inference phase. However, due to the task difference, we extensively modify the editing scheme of BLD for defect generation. First of all, given a defect-free image $\mathbf{x}_{OK}$ and a trimap $\Gamma$ defined before, the corresponding defect mask $\mathbf{M}^{\ast}_{NG}$ is firstly calculated as introduced in Sec. III-B. Then the multi-stage denoising algorithm with content editing is performed and the workflow of this algorithm is illustrated in Alg. 1 where $\odot$ indicates element-wise multiplication and $\neg$ stands for the element-wise “Not” operation on the binary map.
@@
1 $\mathbf{z}_{OK}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}\leftarrow\Omega(\mathbf{x}_{OK})$
@@ 8       $t=t-1$
10/* Latent Editing Stage */
@@
12       $\mathbf{z}_{t}=\mathbf{z}_{t}\odot\mathbf{M}^{z}_{NG}+\mathbf{z}_{OK}\odot\neg\mathbf{M}^{z}_{NG}$ $\mathbf{z}_{t-1}\leftarrow Denoise(\mathbf{z}_{t},\Gamma,y)$
@@ 13       $t=t-1$
15/* Image Editing Stage */
@@ 20       $t=t-1$
return $\mathbf{z}^{\ast}_{NG}=\mathbf{z}_{0}$
@@
As shown in Fig. 2, the pipeline consistes of three stages in the denoising phase. The first stage is the conventional denosing process without content editing and takes $T_{1}$ steps. The second stage takes $T_{2}$ steps and the input feature $\mathbf{z}_{t}$ is merged with the feature of $\mathbf{x}_{OK}$, under the guidance of the dilated defect mask $\mathbf{M}^{\ast}_{NG}$. In the third stage, a novel content blending method proposed in this paper is conducted. In specific, the blending operation is performed in the pixel space rather than the latent space and the frozen encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$ are used to switch the information between the two spaces.
@@ III-E Online Decoder Adaptation
Recall that the aim of our algorithm is to generate better synthetic defective images which are actually reconstructed by the decoder $\Phi(\cdot)$, we thus propose to fine tune the decoder for each generated sample to achieve higher image quality.
Given the blended latent feature $\mathbf{z}^{\ast}_{NG}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ obtained in Alg. 1, the demanding defect mask $\mathbf{M}^{\ast}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}$ and the source image $\mathbf{x}_{OK}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, one can fine tune the decoder model $\Phi(\cdot)$ by conducting the online update algorithm described in Alg. 2
@@
1 $\mathbf{x}^{\ast}_{NG}\leftarrow\Phi(\mathbf{z}^{\ast}_{NG}),~{}\bar{\mathbf{M}}^{\ast}_{NG}(i,j)\leftarrow\neg\mathbf{M}^{\ast}_{NG}(i,j)$
@@ 8       $t=t+1$
One can see that the online adaptation strategy encourages the decoder $\Phi(\cdot)$ generating similar defect-free pixels as $\mathbf{x}_{OK}$ while outputting the defective pixels as the original decoder suggests. The conservative ratio $\lambda_{con}$ balance the two requirements. In the experiments of this work, we prove that the individually updated decoder can yield better samples and lead to higher AD performance.
@@ III-F Implementation Details
In this paper, all the images are resize to $256\times 256$ following the setting of DeSTseg [6] which is a SOTA anomaly detection algorithm and adopted by this paper for evaluating the quality of generated samples. In this way, we set $H_{x}=W_{x}=256$ and $H_{z}=W_{z}=32$. The latent feature channel number $C_{z}$ is $4$ in this paper for efficient denoising operation. As to the trimap prompts, the object’s region is estimated by using the foreground estimation method proposed in [11] and the anomalous mask is obtained based on the randomly selected $10$ genuine defective samples for each subcategory of the dataset. The affine transformation matrix mentioned in Sec. III-B3 is determined by a random rotation angle $\gamma\in[0^{\circ},360^{\circ}]$ and a random scaling factor $s\in[0.85,1.15]$. For the AD tasks on texture subcategories, such as a wooden or a fabric surface, all the pixels belong to the foreground region except the defective ones.
We employ the AdamW optimizer [51] for updating the model parameters, with a default learning rate of $1\times 10^{-5}$, and default betas are set to $0.9$ and $0.999$, respectively. During training, there is a $10\%$ chance that empty text prompt is input to the model.
During the inference stage, we adopt the DDIM [34] sampling method with the normal denoising step $T_{1}=50$, the latent blended step $T_{2}=30$, and pixel blended step $T_{3}=5$. In the online decoder adaptation stage, $T_{ft}=200$ and the conservative ratio $\lambda_{con}=100$ to avoid overfitting to the anomaly-free region. The same optimizer as offline learning is used for this online adaptation except the learning rate is set to $1\times 10^{-4}$.
@@ IV-A6 Training Settings for the Involved AD Models
In this paper, we propose to evaluate the generated samples by using the simplest AD method and a sophisticated AD method. In this way, one can obtain a thorough examination of the generated samples.
As one end of the spectrum, the SVM-based anomaly detection algorithm is extremely simple and thus can directly reflect the patch-wise quality of the generated images. In specific, the SVM model based on the RBF kernel [61] is employed as the patch classifier. The patch features are extracted by using the PatchCore algorithm [4] and each $768$-D feature is assigned a positive label if the corresponding patch is normal while negative labels are assigned to the anomalous ones. In each trial, $100$ images are randomly sampled from the synthetic defective images and then $5000$ positive patch features and $5000$ negative patch features are extracted for training the SVM model. The final performance on each subcategory is calculated by averaging the results of $10$ random trails with the same setting. Besides, the performance of the SVM model trained merely on the genuine defective data is also reported. As to the parameter setting, we by default set the $\gamma$ of the RBF kernel as $1\times 10^{-4}$ while the regularization parameter $C$ is set to $1$.
As a SOTA anomaly detection method, DeSTseg [6] employs a discriminative model to detect anomalous regions based on comparing them and their reconstructed defect-free counterparts. Before training, we first generate $5000$ defective samples by using AdaBLDM and another $5000$ defective samples by using the data augmentation method of DeSTseg. We store the defective samples generated by AdaBLDM in $\mathcal{D}_{gen}$ and store those augmented by DeSTseg in $\mathcal{D}_{raw}$. The training data is then sampled from the defect-free samples, $\mathcal{D}_{raw}$ and $\mathcal{D}_{gen}$ in the ratio of $5:4:1$. Note that according to the methodology of DeSTseg, the genuine defective samples, which are the seeds of the synthetic data generation, can not be used for training as the anomaly-free version of them is unknown. For each subcategory, the DeSTseg is learned for $10000$ iterations and we pick the DeSTseg model with a UNIQUE iteration number for all the subcategories in one dataset.
@@ IV-C Qualitative Results
We further illustrate the comparison in a more fine-grained level in Fig. 5. From the figure we can see that the proposed method is more well-controlled in generating the defect regions with the specific category and location.
@@
In addition, the generated images with synthetic defects on BTAD and KSDD2 are also shown in Fig. 6. One can see that besides MVTec AD, the image quality of AdaBLDM is maintained for the two different datasets.
@@
Finally, as a qualitative complement to Tab. II, Fig. 7 shows the heat map of anomaly predicted by the DeSTseg algorithm. We can see that thanks to the higher quality of the simulated synthetics, the AdaBLDM-equipped AD algorithm can output the best heat maps for anomaly detection and localization.
@@ IV-D Ablation Study
To understand the success of the proposed AdaBLDM algorithm, we utilize ablation experiments to evaluate the specific contributions of the primary components which include:
@@ V Conclusion & future work
In this paper, we propose a sophisticated method to generate realistic defective samples for industrial anomaly detection and localization. Motivated by the great success of the diffusion model, we highly customize the Blended Latent Diffusion Model (BLDM) with several novel modifications for this specific task. The yielded algorithm is termed AdaBLDM and can generate better synthetic samples compared with the existing methods based on conventional “cut-and-paste” scheme or Generative Adversarial Networks (GANs). By applying the AdaBLDM for AD data augmentation, one can observe significant increase on the AD performances of SOTA algorithms. The AIGC algorithm proposed in this paper can be directly used as an effective data augmentation method in real-world applications and also sheds light on the influential factors for solving defect generalization task. In the future, one could obtain even better synthetic defective samples by introducing finer controlling signals such as the language prompts describing the details (shape, size etc.) about the demanding defects. Another appealing direction of future work is to speed up the denoising process of the diffusion model so that more defective samples can be generated given a limited time budget, which is a crucial constraint in most practical scenarios.

### v2

@@ III-B3 Defect trimap prompts
As shown in Fig. 2, in the proposed AdaBLDM algorithm, the controlling trimap $\Gamma$ is firstly embedded by a convolutional block as $\zeta(\Gamma)\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$. The embedding feature is then fed into an encoder network $\hat{\epsilon}_{tri}(\cdot)$ that shares the same structure and initial parameters with the denoising encoder of $\epsilon_{\theta}(\cdot)$, except that the original “spatial attention” module is replaced with a self-attention process over the embedded trimap features. The deep features extracted from different layers of $\hat{\epsilon}_{tri}(\cdot)$ are then injected into the corresponding layers of the denoising decoder of $\epsilon_{\theta}(\cdot)$ via element-wise addition (see Fig. 2).
III-C Loss Function and Training Scheme
Given the cross-modal controlling features $\{\tau(y),\zeta(\Gamma)\}$, the learning objective of the proposed model is defined as:
$L_{LDM}\coloneqq\mathbb{E}_{\mathbf{z}_{t},\mathcal{C},\epsilon\thicksim\mathcal{N}(0,1),t}\left[\|\epsilon-\epsilon_{\theta}(\mathbf{z}_{t},t,\tau(y),\hat{\epsilon}_{tri}(\zeta(\Gamma)))\|^{2}_{2}\right],$
where ${\epsilon}_{\theta}(\cdot)$ denotes the denoising model and $\hat{\epsilon}_{tri}$ stands for the encoder for the latent trimap features. In this work, the involved modules are learned with different training schemes as described below.
  • ${\epsilon}_{\theta}(\cdot)$ is pre-trained by using the domain dataset (e.g., all training images of all the subcategories of MVTec AD [22] for the experiments on this dataset). We keep ${\epsilon}_{\theta}(\cdot)$ frozen during model learning for each subcategory.
  • The trimap feature encoder $\hat{\epsilon}_{tri}(\cdot)$ is initialized by the corresponding parameters of the original LDM model [19] and then fine-tuned for each subcategory based on the real defective samples
  • The convolutional block $\zeta(\cdot)$ for trimap embedding is only learned based on the genuine anomaly samples from scratch.
  • For the image encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$, we follow [19] to deploy the corresponding modules of the VQ-VAE algorithm [53], which are fixed during the training.
  • The text encoder $\tau(\cdot)$ is copied from CLIP and frozen during training.
@@ III-D Multi-Stage Denoising with Content Editing
In this paper, we propose to perform feature editing in the inference phase by tailoring the editing scheme of Blended Latent Diffusion (BLD) [20] for defect generation.
First of all, given a defect-free image $\mathbf{x}_{OK}$ and a trimap $\Gamma$, the corresponding defect mask $\mathbf{M}^{\ast}_{NG}$ is calculated as introduced in Sec. III-B3. Then the multi-stage denoising algorithm with content editing is performed and the workflow of this algorithm is illustrated in Alg. 1 where $\odot$ indicates element-wise multiplication and $\neg$ stands for the element-wise operation of logical “NOT” on the binary map.
1
@@
3
[…]
@@ 12       $t=t-1$
14/* Latent Editing Stage */
@@
16
17      $\mathbf{z}_{t}=\mathbf{z}_{t}\odot\mathbf{M}^{z}_{NG}+\mathbf{z}_{OK}\odot\neg\mathbf{M}^{z}_{NG}$
18      $\mathbf{z}_{t-1}\leftarrow Denoise(\mathbf{z}_{t},\Gamma,y)$
@@ 19       $t=t-1$
21/* Image Editing Stage */
@@ 26       $t=t-1$
28
29return $\mathbf{z}^{\ast}_{NG}=\mathbf{z}_{0}$
@@
As shown in Fig. 3, the pipeline consists of three stages in the denoising phase. The first stage takes $T_{1}$ steps to conduct the conventional denosing[38, 39, 19] without content editing. The second stage spends $T_{2}$ steps and the input feature $\mathbf{z}_{t}$ is merged with the feature of $\mathbf{x}_{OK}$, under the guidance of the dilated defect mask $\mathbf{M}^{\ast}_{NG}$. In the third stage, a novel content blending method proposed in this paper is conducted. In specific, the blending operation is performed in the pixel space and the blended image is then remapped back to the latent space for denoising. The image encoder $\Omega(\cdot)$ and decoder $\Phi(\cdot)$ are used cooperatively to switch the information between the two spaces.
@@ III-E Online Decoder Adaptation
Considering that the image quality of the synthetic defective sample can be significantly influenced by the image decoder $\Phi(\cdot)$, we propose to fine-tune the decoder for each generated sample. Once the blended latent feature $\mathbf{z}^{\ast}_{NG}\in\mathbb{R}^{H_{z}\times W_{z}\times C_{z}}$ is obtained from Alg. 1, and given the defective mask $\mathbf{M}^{\ast}_{NG}\in\mathbb{B}^{H_{x}\times W_{x}}$ and the source image $\mathbf{x}_{OK}\in\mathbb{R}^{H_{x}\times W_{x}\times 3}$, we can fine-tune the decoder model $\Phi(\cdot)$ by conducting the online update algorithm as described in Alg. 2, which shows that the online adaptation strategy encourages the decoder $\Phi(\cdot)$ generating defect-free pixels assembling to $\mathbf{x}_{OK}$ and simultaneously outputting the defective pixels guided by the original decoder. The conservative ratio $\lambda_{con}$ balances the two objectives. In the experiments of this paper, we demonstrate that this online-updated decoder can generate more realistic samples for AD training and thus leads to higher AD performance.
1
@@
3
[…]
@@ III-F Implementation Details
In this paper, all the images are resized to $256\times 256$ following the setting of DeSTseg [6] which is a SOTA anomaly detection algorithm adopted by this paper for evaluating the effectiveness of the generated samples. In this way, we set $H_{x}=W_{x}=256$ and $H_{z}=W_{z}=32$. The latent feature channel number $C_{z}$ is $4$ in this paper for efficient denoising operation. As to the trimap prompts, the object’s region is estimated by using the foreground estimation method proposed in [14] and the anomalous mask is obtained based on the randomly selected $10$ genuine defective samples for each subcategory of the dataset. The affine transformation matrix mentioned in Sec. III-B3 is determined by a random rotation angle $\gamma\in[0^{\circ},360^{\circ}]$ and a random scaling factor $s\in[0.85,1.15]$. For the AD tasks on texture subcategories, such as a wooden or a fabric surface, all the pixels are considered as foreground except the defective ones.
We employed the AdamW optimizer [54] for updating the model parameters, with a default learning rate of $1\times 10^{-5}$, and default betas are set to $0.9$ and $0.999$, respectively. During training, there is a $10\%$ chance that an empty text prompt is input to the model. During the inference stage, we adopted the DDIM [39] sampling method with the normal denoising step $T_{1}=50$, the latent blended step $T_{2}=30$, and pixel blended step $T_{3}=5$. In the online decoder adaptation stage, $T_{ft}=200$ and the conservative ratio $\lambda_{con}=100$ to avoid overfitting to the anomaly-free region. The same optimizer as offline learning is used for this online adaptation except for the learning rate is set to $1\times 10^{-4}$.
@@ IV-A6 Training Settings for the Involved AD Models
In our experiments, the generated samples are evaluated by using the simplest AD method and a sophisticated AD method. As one end of the spectrum, the SVM-based anomaly detection algorithm is extremely simple and thus can directly reflect the primitive quality of the generated images, at the patch level. In specific, the SVM model based on the RBF kernel [64] is employed as the patch classifier. The patch features are extracted by using the PatchCore algorithm [4] and each $768$-D feature is assigned a positive label if the corresponding patch is normal while negative labels are assigned to the anomalous ones. In each trial, $100$ images are randomly sampled from the synthetic defective images and then $5000$ positive patch features and $5000$ negative patch features are extracted for training the SVM model. The final performance on each subcategory is calculated by averaging the results of $10$ random trails with the same setting. Besides, the performance of the SVM model trained merely on the genuine defective data is also reported. As to the parameter setting, we by default set the $\gamma$ of the RBF kernel as $1\times 10^{-4}$ while the regularization parameter $C$ is set to $1$.
Referring to DeSTseg [6] which employs a discriminative model to detect anomalous regions by comparing them with reconstructed defect-free counterparts, before training, we first generate $5000$ defective samples by using AdaBLDM and another $5000$ defective samples by using the data augmentation method of DeSTseg. We store the defective samples generated by AdaBLDM in $\mathcal{D}_{gen}$ and store those augmented by DeSTseg in $\mathcal{D}_{raw}$. The training data is then sampled from the defect-free samples, $\mathcal{D}_{raw}$ and $\mathcal{D}_{gen}$ in the ratio of $5:4:1$. Note that according to the methodology of DeSTseg, the genuine defective samples, which are the seeds of the synthetic data generation, can not be used for training as the anomaly-free version of them is unknown. For each subcategory, the DeSTseg is learned for $10000$ iterations and we pick the DeSTseg model with a UNIQUE iteration number for all the subcategories in one dataset.
@@ IV-C Qualitative Results
We further compare with existing methods in fine-grained details, and experimental results are reported in Fig. 5. As shown in Fig. 5, the proposed method is more well-controlled in generating the defect regions with the specific category and location. In addition, the generated images with synthetic defects on BTAD and KSDD2 are shown in Fig. 6. It can be seen that besides MVTec AD, the image quality of AdaBLDM is maintained for the two different datasets.
@@
Finally, as a qualitative complement to Tab. II, Fig. 7 shows the heat map of anomaly predicted by the DeSTseg algorithm. We can see that high-quality simulated synthetics benefits the AdaBLDM-equipped algorithm which can output competitive heat maps for anomaly detection and localization.
@@ IV-D Ablation Study
In this section, we systematically evaluate the contribution of each component in our method. The following experiments are conducted:
@@ V Conclusion & future work
In this paper, we propose an advanced method to generate realistic defective samples for industrial anomaly detection and localization. Inspired by the diffusion model, we customize the Blended Latent Diffusion Model (BLDM) to be well-suited to this specific task. The proposed algorithm, termed AdaBLDM, can generate better synthetic samples compared with existing methods based on conventional “cut-and-paste” schemes or Generative Adversarial Networks (GANs). By applying the AdaBLDM in anomaly detection data augmentation, we achieve a state-of-the-art performance of AD, thanks to the well-controlled defective regions aligned at the pixel level. In the future, we explore high-quality synthetic defective samples by introducing finer controlling signals such as the language prompts describing the details (shape, size, etc.) of the demanding defects. Another appealing direction of future work is to speed up the denoising process of the diffusion model so that more defective samples can be generated given a limited time budget, which is a crucial constraint in most practical scenarios.